OpenStudy (anonymous):

find the exact value cos -105°

OpenStudy (anonymous):

I am guessing that you will have to use the sum formula: \[\cos(x+y) = \cos(x)*\cos(y) - \sin(x)*\sin(y)\]

OpenStudy (anonymous):

or difference formula.

OpenStudy (anonymous):

-0.24095904924

OpenStudy (anonymous):

that's the exact value

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=cos-105

OpenStudy (anonymous):

That is NOT the exact value.

OpenStudy (anonymous):

105=60+45 Use @EulersEquation's formula. \[\cos (-105 {}^{\circ})=-\frac{\sqrt{3}-1}{2 \sqrt{2}} \]